Project

General

Profile

01062014 Minutes

Attending: Mat, Gavin, Michael

-Temperature at Fermilab -15F.
-Operations: APD water leak. One APD on dcm-2-02-06 replaced on Saturday.
-Andrew had start send the data to enstore (dCache). That system is being worked out.
-Only 3 TB of disk.

-Trying to apply masks to and bad channels to existing NUmI
-Some dB issues.

-Database connection issues - Michael has had repeated issues as well.

Send Collab session agenda around:
Would like the following to attend: Steve, Jon, Peter …